OpenStudy (dls):

The sums of n terms of two AP's are in the ratio 5n+4:9n+6..Find ratio of 18th terms

OpenStudy (dls):

\[\LARGE \frac{a+(n-1)d}{a'+(n-1)d'}=\frac{5n+4}{9n+6}\]

OpenStudy (dls):

I heard we do some hit n trial here,idk that

OpenStudy (shubhamsrg):

not hit and trial, we just HIT!

OpenStudy (dls):

:|

OpenStudy (shubhamsrg):

sum of n terms = S_n ,(say) nth term = a_n, (say) Can you say that S_1 = a_1 ?

OpenStudy (dls):

yes

OpenStudy (shubhamsrg):

So you can find the ratio of 1st terms of both the APs now ?

OpenStudy (dls):

no:P

OpenStudy (shubhamsrg):

just substitute n=1

OpenStudy (dls):

:O

OpenStudy (anonymous):

Can you substitute n=18 ?

OpenStudy (shubhamsrg):

better substitute n=35! :P